OpenStudy (anonymous):

-3÷2i in standard form

OpenStudy (anonymous):

its easy just mulitiply numerator and denominator by i you get -3i/2i^2 =-3i/-2 as we know i^2=-1 =3i/2

OpenStudy (anonymous):

\[\frac{ -3 }{ 2\iota },multiply numerator and denominator by \]